OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4

OpenStudy (anonymous):

Answers are either A: 2 B: 3 C: All Real Numbers D: No Solution

OpenStudy (anonymous):

−2(x + 3) = −2(x + 1) − 4 Order of opperations: Multiply -2(x+3) 's first

OpenStudy (anonymous):

-2x-2=-2x-2 no solution

OpenStudy (anonymous):

Thanks :)